OpenStudy (anonymous):

solve for h, V=3.14r^2h

OpenStudy (hba):

\[V= \pi r^2h\] divide whole equation by \[\pi r^2\]

OpenStudy (anonymous):

\[\pi r ^{2}\] is. would it just be \[ \frac{ v }{ \pi r^2 }\] ?

OpenStudy (hba):

So \[h=\frac{ V }{ \pi r^2 }\] Yeah :D

OpenStudy (anonymous):

thanks.

OpenStudy (hba):

Welcome @mom22